Palencia Paints Corporation has a target capital structure of 45% debt and 55% common equity, with no preferred stock. Its before-tax cost of debt is 12% and its marginal tax rate is 40%. The current stock price is P0 = $25.00. The last dividend was D0 = $2.25, and it is expected to grow at a 4% constant rate. What is its cost of common equity and its WACC? Round your answers to two decimal places. Do not round your intermediate calculations.

  1. rs =  %

  2. WACC =  %
Answer #1

cost of common equity=(D1/Current price)+Growth rate

=[(2.25*1.04)/25]+0.04

=13.36%

Cost of debt after-tax=12*(1-tax rate)

=12*(1-0.4)=7.2%

WACC=Respective costs*Respective weight

=(13.36*0.55)+(0.45*7.2)

=10.59%(Approx).